A girls mother is 28 years older than she. In 4 years her mother will be 3 times her age. How old is the mother now?
I tried:
d = daughter
m = mother
d = m - 28
m = d + 28
d + 4 = 1/3(m + 4)

Answer by scott8148(6628)

m = g + 28 ___ m - 28 = g

m + 4 = 3(g + 4) ___ m + 4 = 3g + 12

substituting ___ m + 4 = 3(m - 28) + 12 = 3m - 84 + 12

76 = 2m ___ 38 = m